OpenStudy (anonymous):

(m-9)+6=4(m-2)-2(m-6)

OpenStudy (anonymous):

is there a question with it

OpenStudy (anonymous):

it just says solve the equation

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

O RLY ?

OpenStudy (anonymous):

yup hold on

OpenStudy (anonymous):

1 Simplify brackets m−9+6=4(m−2)−2(m−6) 2 Simplify m−9+6 to m−3 m−3=4(m−2)−2(m−6) 3 Expand m−3=4m−8−2m+12 4 Simplify 4m−8−2m+12 to 4m−2m+4 m−3=4m−2m+4 5 Simplify 4m−2m+4 to 2m+4 m−3=2m+4 6 Subtract m from both sides −3=2m+4−m 7 Simplify 2m+4−m to m+4 −3=m+4 8 Subtract 4 from both sides −3−4=m 9 Simplify −3−4 to −7 −7=m 10 Switch sides m=−7

OpenStudy (anonymous):

wtf is with all the question marks

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

wait for it to load

OpenStudy (anonymous):

say whallhi thank you